# On optimal solutions of classical and sliced Wasserstein GANs with non-Gaussian data

### v1

@@ IV-C Proof of Lemma 2
The R.H.S of (28) can be solved by following the proof steps in Sec. IV-A. When $q=2$, for any $\omega$,
$\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]$ (58)
$=$ $\int_{\mathbb{R}}|\omega^{T}x-t^{\Theta}_{\omega}(\omega^{T}x)|^{2}d\mu_{\omega}(\omega^{T}x),$
where the optimal trasport function
@@
For the L.H.S of (28), when $q=2$, for any $\omega$,
$\inf_{\pi\in\Pi(\mu,\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]$ (60)
$=$ $\int_{\mathbb{R}^{d}}|\omega^{T}x-t^{\Theta}(x)|^{2}d\mu(x),$ (61)
where the optimal transport function $t^{\Theta}(x)$ equals to $t^{\Theta}_{\omega}(\omega^{T}x)$ in (59), which will be proved in Lemma 8 later. Then, with the change of variables[28], (61) and (58) are equal.
To show that the optimal transport function $t^{\Theta}(x)$ for (60) equals to $t^{\Theta}_{\omega}(\omega^{T}x)$ in (59), we first prove that (60) belongs to the general multi-to-one optimal transport problem in [21] by checking its properties. To do this, let $Y=\omega^{T}G_{\Theta}(Z)$ and introduce the following sets. For each fixed $y\in\mathbb{R}$, for all $k\in\mathbb{R}$, we define sets
$X(y,k)$ $:=\left\{x\in\mathbb{R}^{d}:-\frac{\partial}{\partial y}c(x,y)=k\right\}$
$=\left\{x\in\mathbb{R}^{d}:\omega^{T}x=y+{k}/{2}\right\},$
$X_{\leq}(y,k)$ $:=\left\{x\in\mathbb{R}^{d}:-\frac{\partial}{\partial y}c(x,y)\leq k\right\}$
$=\left\{x\in\mathbb{R}^{d}:\omega^{T}x\leq y+{k}/{2}\right\},$ (62)
@@
(a) $c(x,y)$ is non-degenerate in the sense of [21, Definition 3.1] since the rank of $D^{2}_{xy}c(x,y)\geq 1$ for all $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}$. The desired result follows from [21, Theorem 4(a)].
$\frac{\partial^{2}}{\partial x\partial y}c^{w}(x,y)=\begin{bmatrix}2ww^{T}&-2w\\ -2w^{T}&2\end{bmatrix}$.
@@
$X_{\leq}(y,k^{+}(y))=\{x\in\mathbb{R}^{d}:w^{T}x\leq f^{+}(y)\}$ (65)
$\supseteq$ $\{x\in\mathbb{R}^{d}:w^{T}x\leq f^{+}(y^{\prime})\}=X_{\leq}(y^{\prime},k^{+}(y^{\prime})),$
@@
$\mu(X_{\leq}(y,k^{+}(y)))=\nu^{\Theta}_{w}((-\infty,y))$ (66)
$\leq$ $\nu^{\Theta}_{w}((-\infty,y^{\prime}))=\mu(X_{\leq}(y^{\prime},k^{+}(y^{\prime}))).$
@@
$\mu(X_{\leq}(y^{\prime},\bar{k}))=\mu(X_{\leq}(y,k^{+}(y)))$
$=\nu^{\Theta}_{w}((-\infty,y))=\nu^{\Theta}_{w}((-\infty,y^{\prime})),$
@@
Then, the desired result follows from [21, Theorem 4(b)]. ∎
@@
V Empirical study on convergence with synthetic data
@@
We have only considered population WGAN thus far. In practice, the distribution is estimated from training data and does not have a closed-form CDF. Therefore, it is of interest to empirically study how fast the solution converges to the population WGAN result using synthetic training data. In Fig. 1, we consider the linear activation and plot the optimal $\theta_{2}^{*}$ in (16)(17) obtained by solving population WGAN with $d=1,q=2$, and its estimate with synthetic data when $\mu$ is chosen to be (a) $\mathcal{N}(0,1)$ and (b) Laplace distribution [29] with mean 0 and scale $1/\sqrt{2}$ (which also has a unit variance). Specifically, by generating training data $\{x_{i}\}^{M}_{i=1}$ according to the true distribution $\mu$, we empirically estimate $\theta_{2}^{*}$ from (16) by
$\frac{1}{M}\sum^{M}_{i=1}x_{i}\Phi^{-1}(\hat{F}_{\mu}(x_{i}))$ (71)
@@
and the kernel density estimation (KDE) [30] is used to replace the true CDF $F_{\mu}(x)$ with the estimated one $\hat{F}_{\mu}(x)$ from the training data. For each distribution, our result shows a nice convergence behavior where the difference becomes smaller than 0.005 with only MM=50000 data size. The optimal $\theta_{2}^{*}$ in (16)(17) is 1 and 0.98013 for Gaussian and Laplace distributions respectively. Using linear generator, GAN output is also Gaussian and thus the convergence behavior for Gaussian data is better than that for Laplace data. Note that $\theta^{*}_{1}$ in (16) can be estimated by the sample mean, so the convergence is not shown in Fig. 1. We also use SGD with momentum to estimate $\theta_{1}$ and $\theta_{2}$ for loss function (1) under $d=1,q=1$, with gradient empirically obtained from (18) using KDE $\hat{F}_{\mu}(x)$. With $\mathcal{N}(\mu=1.5,\sigma^{2}=4)$ data, $(\theta_{1},\theta_{2})$ converges to $(1.4957,2.0905)$ shown in Fig. 2 and close to the population local optimum $(\mu,\sigma)$ for $q=1$.
The convergence rate of the empirical WGAN solution to that of the population WGAN problem has been theoretically analyzed to be $M^{-2/d}$ for the LQG setting in [8, Theorem 2]. Note that the convergence of iteratively solving empirical WGAN based on a good solver for (2) was given in [12]. However, this convergence heavily relies on a good semi-discrete optimal transport solver, which is still hard to design [31]. It is our future work to analyze the convergence rate for our non-Gaussian results, also compare it with the empirical simulations for $d>1$.
@@
Figure 1: Comparison of optimal parameters in (16)(17) with their estimates (71) using synthetic data.
Figure 2: the estimation of $\theta_{1}$ and $\theta_{2}$ in W1Gaussain via SGD
@@
Figure 3: Comparison of sliced Wasserstein distance with our optimal parameter and [8]’s when data is i.i.d.
@@
Figure 4: Comparison of sliced Wasserstein distance with our optimal parameter and [8]’s when data is correlated.
@@
For the sliced WGAN problem, we compare our asymptotically optimal linear generator in theorem 3 with the $r$-PCA solution which is the optimal WGAN solution in the LQG setting. First, We use synthetic data $\{X_{1:d}^{(i)}\}_{i=1}^{M}$ with data size $M=4\times 10^{5}$, where $\{X_{1:d}^{(i)}\}$ generated follows i.i.d Laplace distribution with zero mean and unit variance for all $i$, and estimate the empirical distribution $\mu=M^{-1}\sum_{i=1}^{M}\delta_{X^{(i)}}$. We compute the empirical sliced Wasserstein distance with Monte Carlo based on $2\times 10^{4}$ random projections,
$\left(\sum_{\omega_{i}\in\Omega,\ i=1}^{20000}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]d\omega\right)^{\frac{1}{2}}.$ (72)
@@
Then, we let $\Theta$ of (72) be our proposed linear generator parameters (19) and the parameters of $r$-PCA solution. In Fig. 3, we can see that the sliced Wasserstein distance with our proposed linear generator parameters is much lower than the one of $r$-PCA when $r=\frac{2}{3}d$.
In addition, we use not only the i.i.d. synthetic data but also the data generated by AR model mentioned in Remark 1. The AR model is defined as $X_{j}=0.5X_{j-1}+\eta_{j}$ where $\eta_{j}$ is i.i.d. student distribution noise. Given the correlated data, we compare the value of (72) with our proposed linear generator parameters to the value with the parameters of $r$-PCA. In Fig. 4, except for $d=10$, we can see that the sliced Wasserstein distance with our proposed linear generator parameters is lower like the i.i.d. data scenario when $r=\frac{2}{3}d$.
@@
When $r=d$, our results reach the same value as $r$-PCA method with both i.i.d. data and correlated data at $d=1000$ and $d=2000$, respectively. Although $r$-PCA method can have lower sliced Wasserstein distance in some case like $r=d$, the computation complexity of solving the $r$-PCA optimizer is higher than the complexity of our asympotically optimizer. For the $r$-PCA optimizer, the complexity of computing the covariance and find the eigenvalues is $O((M+r)d^{2})$. For our proposed linear generator parameters, we only need to compute $\mathbb{E}_{\mu}[\|X\|^{2}]$ which complexity is $O(Md)$. Our asymptotically optimal linear generator has comparable performance to the r-PCA solution while reducing the computation energy resources.
@@ Acknowledgement
Appendix A Proof for Theorem 2
As $\Psi$ is neither continuous nor strictly increasing, (31) does not hold in general and we cannot directly apply proofs in Sec. IV-A. Again assume $\theta_{2}\in\mathbb{R}_{+}$, Under the condition $\theta_{2}\geq 0$, we deduce from (29) and (12) that
Similarly to (35), the sub-problem (33) now becomes
$\begin{split}&\min_{\theta_{1},\theta_{2}\in\mathbb{R}}J(\theta_{1},\theta_{2}):=\int_{\{F_{\mu}(x)\leq 1/2\}}\Big{(}\theta_{1}-x\Big{)}^{2}F_{\mu}^{\prime}(x)dx\\ &\hskip 21.68121pt+\int_{\{F_{\mu}(x)>1/2\}}\Big{(}\theta_{1}+\theta_{2}\Phi^{-1}(F_{\mu}(x))-x\Big{)}^{2}F_{\mu}^{\prime}(x)dx\\ &\hbox{subject to}\quad g(\theta_{1},\theta_{2}):=-\theta_{2}\leq 0.\end{split}$ (74)
@@ Appendix A Proof for Theorem 2
$\int_{\{F_{\mu}(x)\leq 1/2\}}\Big{(}\theta_{1}-x\Big{)}F_{\mu}^{\prime}(x)dx$ (75)
$+$ $\int_{\{F_{\mu}(x)>1/2\}}\Big{(}\theta_{1}+\theta_{2}\Phi^{-1}(F_{\mu}(x))-x\Big{)}F_{\mu}^{\prime}(x)dx=0,$
$2\int_{\{F_{\mu}(x)>1/2\}}\Big{(}\theta_{1}+\theta_{2}\Phi^{-1}(F_{\mu}(x))-x\Big{)}$ (76)
$\Phi^{-1}(F_{\mu}(x))F_{\mu}^{\prime}(x)dx-\lambda=0.$
Let $X$ be a random variable with cdf $F_{\mu}$. Then, (75)-(76) can be written as
$\theta_{1}\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]$ (78)
$+$ $\theta_{2}\mathbb{E}\left[(\Phi^{-1}(F_{\mu}(X)))^{2}1_{\{F_{\mu}(X)>1/2\}}\right]$
- $\mathbb{E}\left[X\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]-\lambda/2=0.$
Plugging $\theta_{1}=\mathbb{E}[X]-\theta_{2}\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]$ (from (77)) into (78), we obtain
$\mathbb{E}[X]\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]$
$+$ $\theta_{2}\text{Var}\left(\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)$
- $\mathbb{E}\left[X\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]-\lambda/2=0.$ (79)
Hence, a solution $(\theta_{1},\theta_{2},\lambda)$ to (36) must equivalently satisfy (77) and (79). Note that $F_{\mu}(X)\sim\operatorname{Uniform}[0,1]$, so that the cdf of $\Phi^{-1}(F_{\mu}(X))$ is simply $\Phi$. That is, $\Phi^{-1}(F_{\mu}(X))$ and $Z$ have identical distribution. It follows that
$\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]$
$=$ $\mathbb{E}[Z1_{\{\Phi(Z)>1/2\}}]=\mathbb{E}[Z1_{\{Z>0\}}]=\frac{1}{\sqrt{2\pi}},$ (80)
$\mathbb{E}\left[(\Phi^{-1}(F_{\mu}(X)))^{2}1_{\{F_{\mu}(X)>1/2\}}\right]$
$=$ $\mathbb{E}[Z^{2}1_{\{\Phi(Z)>1/2\}}]=\mathbb{E}[Z^{2}1_{\{Z>0\}}]=\frac{1}{2}.$ (81)
In view of this, (77) and (79) reduce to
@@
As $\Phi^{-1}$ is strictly increasing and $F_{\mu}$ is nondecreasing, the map $x\mapsto\Phi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}$ is nondecreasing. Hence, by a similar argument in (40), $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)\geq 0$. The result then follows by repeating the same arguments in the proof of Lemma 1. Specifically, we deal with the two cases where $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)$ is strictly positive and zero separately. In each case, we solve (82) (in place of (38) in Lemma 1) and find that $H_{J}+\lambda H_{g}$ now takes the form
$\begin{bmatrix}1&\frac{1}{\sqrt{2\pi}}\\ \frac{1}{\sqrt{2\pi}}&\frac{1}{2}\end{bmatrix}$,
which is positive definite. ∎
@@
$\begin{split}&\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)\\ =&\operatorname{Var}(X)-\frac{2\pi}{\pi-1}\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)^{2}.\end{split}$ (84)

### v2

@@ IV-C Proof of Lemma 2
The R.H.S. of (29) can be solved by following the proof steps in Sec. A. When $q=2$, for any $\omega$,
$\inf_{\pi_{\omega}\in\Pi_{\omega}(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi_{\omega}}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]=\int_{\mathbb{R}}|\omega^{T}x-t^{\Theta}_{\omega}(\omega^{T}x)|^{2}d\mu_{\omega}(\omega^{T}x),$ (70)
where the optimal transport function
@@
For the L.H.S. of (29), when $q=2$, for any $\omega$,
$\inf_{\pi^{\prime}\in\Pi^{\prime}(\mu,\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi^{\prime}}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]=\int_{\mathbb{R}^{d}}|\omega^{T}x-t^{\Theta}(x)|^{2}d\mu(x),$ (72)
where the optimal transport function $t^{\Theta}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ has the same value as $t^{\Theta}_{\omega}(\omega^{T}x)$ in (71) given $\omega$, which will be proved in Lemma 9 later. Then, with the change of variables[6], (72)and (70) are equal.
To show that the optimal transport function $t^{\Theta}(x)$ for (72) equals $t^{\Theta}_{\omega}(\omega^{T}x)$ in (71), we first prove that (72) belongs to the general multi-to-one optimal transport problem in [9] by checking its properties. To do this, let $Y=\omega^{T}G_{\Theta}(Z)$ and introduce the following sets. For each fixed $y\in\mathbb{R}$, for all $k\in\mathbb{R}$, we define sets
$X(y,k)$ $:=\left\{x\in\mathbb{R}^{d}:-\frac{\partial}{\partial y}c(x,y)=k\right\}=\left\{x\in\mathbb{R}^{d}:\omega^{T}x=y+{k}/{2}\right\},$
$X_{\leq}(y,k)$ $:=\left\{x\in\mathbb{R}^{d}:-\frac{\partial}{\partial y}c(x,y)\leq k\right\}=\left\{x\in\mathbb{R}^{d}:\omega^{T}x\leq y+{k}/{2}\right\},$ (73)
@@
(a) $c(x,y)$ is non-degenerate in the sense of [9, Definition 3.1] since the rank of $H_{c}\geq 1$ for all $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}$, where $H_{c}$ is the Hessian matrix of $c(x,y)$. The desired result follows from [9, Theorem 4(a)].
$H_{c}=\begin{bmatrix}2ww^{T}&-2w\\ -2w^{T}&2\end{bmatrix}$.
@@
$X_{\leq}(y,k^{+}(y))=\{x\in\mathbb{R}^{d}:w^{T}x\leq f^{+}(y)\}\supseteq\{x\in\mathbb{R}^{d}:w^{T}x\leq f^{+}(y^{\prime})\}=X_{\leq}(y^{\prime},k^{+}(y^{\prime})),$ (76)
@@
$\mu(X_{\leq}(y,k^{+}(y)))=\nu^{\Theta}_{w}((-\infty,y))\leq\nu^{\Theta}_{w}((-\infty,y^{\prime}))=\mu(X_{\leq}(y^{\prime},k^{+}(y^{\prime}))).$ (77)
@@
$\mu(X_{\leq}(y^{\prime},\bar{k}))=\mu(X_{\leq}(y,k^{+}(y)))=\nu^{\Theta}_{w}((-\infty,y))=\nu^{\Theta}_{w}((-\infty,y^{\prime})),$
@@
Then, the desired result follows from [9, Theorem 4(b)]. ∎
@@
V Empirical study with synthetic data
@@
We have only considered population WGAN thus far. In practice, the distribution is estimated from training data and does not have a closed-form CDF. Therefore, it is of interest to empirically study the performance of the population WGAN result using synthetic training data. In Fig. 1, we consider the linear activation and plot the optimal $\theta_{2}^{*}$ in (17)(18) obtained by solving population WGAN with $d=1,q=2$, and its estimate with synthetic data when $\mu$ is chosen to be (a) $\mathcal{N}(0,1)$ and (b) Laplace distribution [4] with mean 0 and scale $1/\sqrt{2}$ (which also has a unit variance). Specifically, by generating training data $\{x^{(i)}\}^{M}_{i=1}$ according to the true distribution $\mu$, we empirically estimate $\theta_{2}^{*}$ from (17) by
$\frac{1}{M}\sum^{M}_{i=1}x^{(i)}\Phi^{-1}(\hat{F}_{\mu}(x^{(i)}))$ (82)
@@
and the kernel density estimation (KDE) [19] is used to replace the true CDF $F_{\mu}(x)$ with the estimated one $\hat{F}_{\mu}(x)$ from the training data. Details are provided in the Appendix F. For each distribution, our result shows a nice convergence behavior where the difference becomes smaller than 0.005 with only MM=50000 sample size. The optimal $\theta_{2}^{*}$ in (17)(18) is 1 and 0.98013 for Gaussian and Laplace distributions, respectively. Using a linear generator, the GAN output is also Gaussian, and thus the convergence behavior for Gaussian data is better than that for Laplace data. Note that $\theta^{*}_{1}$ in (17) can be estimated by the sample mean, so the convergence is not shown in Fig. 1. We also use SGD with momentum to estimate $\theta_{1}$ and $\theta_{2}$ for loss function (1) under $d=1,q=1$, with gradient empirically obtained from (19) using KDE $\hat{F}_{\mu}(x)$. With $\mathcal{N}(\mu=1.5,\sigma^{2}=4)$ data, $(\theta_{1},\theta_{2})$ converges to $(1.4957,2.0905)$ shown in Fig. 2 and close to the population local optimum $(\mu,\sigma)$ for $q=1$. The convergence rate of the empirical WGAN solution to that of the population WGAN problem has been theoretically analyzed to be $M^{-2/d}$ for the LQG setting in [12, Theorem 2]. Note that the convergence of iteratively solving empirical WGAN based on a good solver for (2) was given in [8]. However, this convergence heavily relies on a good semi-discrete optimal transport solver, which is still hard to design empirically[32].
@@
Figure 1: Comparison of optimal parameters in (17)(18) with their estimates (82) using synthetic data under $q=2,d=1$
Figure 2: Convergence of $\theta_{1}$ and $\theta_{2}$ form SGD using (1) for $q=1,d=1$
@@
Figure 3: Comparison of sliced Wasserstein distance with our linear generator (20) and that from r-PCA [12] when data is i.i.d. Laplace
@@
Figure 4: Comparison of sliced Wasserstein distance with our linear generator (20) and that from r-PCA [12] when data is correlated AR model
@@
For high-dimensional sliced WGAN with $d>1$, we compare our asymptotically optimal linear generator in Theorem 2 with the $r$-PCA solution. The r-PCA is widely adopted [13], and is the optimal solution of classical population WGAN in the LQG setting [12]. First, we use synthetic data $\{X^{(i)}\}_{i=1}^{M}$ and $\{Z^{(i)}\}_{i=1}^{M}$ with sample size $M=4\times 10^{5}$, where $X^{(i)}$ generated from i.i.d Laplace distribution with zero mean and unit variance for all $i$ as [25], and $Z^{(i)}$ generated from i.i.d Gaussian distribution with zero mean and unit variance for all $i$. To solve the infimum of (6) in empirical, we sort $\{\omega_{i}^{T}X^{(j)}\}_{j=1}^{M}$ and $\{\omega_{i}^{T}\Theta Z^{(j)}\}_{j=1}^{M}$ as $\omega_{i}^{T}X]_{(1)}\leq[\omega_{i}^{T}X]_{(2)}\leq\cdots\leq[\omega_{i}^{T}X]_{(M)}$ and $\omega_{i}^{T}\Theta Z]_{(1)}\leq[\omega_{i}^{T}\Theta Z]_{(2)}\leq\cdots\leq[\omega_{i}^{T}\Theta Z]_{(M)}$ for each $i$. We compute the empirical sliced Wasserstein distance with Monte Carlo simulation based on $M_{\omega}=2\times 10^{4}$ random projection samples from the unit sphere (4),
$\left(\frac{1}{M_{\omega}}\frac{1}{M}\sum_{\omega_{i}\in\Omega,\ i=1}^{M_{\omega}}\sum_{j=1}^{M}|[\omega_{i}^{T}X]_{(j)}-[\omega_{i}^{T}\Theta Z]_{(j)}|^{2}\right)^{\frac{1}{2}}.$ (83)
@@
We let $\Theta$ of (83) be our proposed linear generator parameters (20) and the parameters of the $r$-PCA solution. In Fig. 3, we can see that the sliced Wasserstein distance with our linear generator is much lower than that of $r$-PCA when $r=\frac{2}{3}d$.
@@
When $r=d$, our results reach the same values as the $r$-PCA method at $d=1000$. The computational complexity of solving our proposed asymptotically optimal parameters is lower than the complexity of the $r$-PCA optimal parameters. For the $r$-PCA optimal parameters, the complexity of computing the covariance and finding the eigenvalues is $O((M+r)d^{2})$, which is $O(Md^{3})$ in the settings $r=\frac{2}{3}d$ and $r=d$. For our proposed linear generator parameters, we only need to compute $\mathbb{E}_{\mu}[\|X\|^{2}]$ whose complexity is $O(Md)$. Our asymptotically optimal linear generator has comparable performance to the r-PCA solution while reducing the computational resources required for parameters.
As [25], in addition to the i.i.d. synthetic data, we also use the data generated from the AR model in Remark 1. For each sample $X^{(i)}$, the $j$-th element $X^{(i)}_{j}=0.5X^{(i)}_{j-1}+\eta_{j}$ where $\eta_{j}$ is i.i.d. student’s-t distribution noise. Given the correlated data, we compare the value of (83) with our proposed linear generator parameters to that of $r$-PCA. In Fig. 4, except for very small $d=10$, we can see that the sliced Wasserstein distance with our proposed linear generator is lower than the i.i.d. data scenario when $r=\frac{2}{3}d$. When $r=d$, our results reach the same values as the $r$-PCA method at $d=2000$ with lower complexity than in the i.i.d. generated data.
VI Conclusion
In this work, we investigated the optimal solution characterization of population WGAN beyond the LQG setting. We first derived closed-form optimal parameters for one-dimensional WGAN when the NN has non-linear activation functions, and the data is non-Gaussian. For high‐dimensional data, we adopted the sliced Wasserstein framework and proved that linear generators can be asymptotically optimal. We further proposed an unprojected sliced WGAN that constrains the whole data marginal rather than the projected one in the original sliced WGAN, and we identified its asymptotic optimality as well. Empirical studies have shown that our generator for sliced WGAN can achieve better performance with linear complexity in the data dimension, compared to the celebrated r-PCA solution, which has cubic complexity.
@@ Acknowledgement
Appendix A Proof for Proposition 1
If $\Psi$ is continuous and strictly increasing and $\theta_{2}\in\mathbb{R}_{+}$, (30) can be expressed in closed-form as
$t^{\Theta}(x)=\theta_{1}+\theta_{2}\Psi^{-1}(F_{\mu}(x)).$ (84)
To see this, since $\theta_{2}>0$, observe that
where the last equality follows from (30) and that $F_{\nu^{\Theta}}$ is continuous and strictly increasing; here, $F_{\nu^{\Theta}}$ inherits the same properties from $\Psi$ thanks to (8). Though the continuity of $\Psi$ is not needed in Lemma 3 and its Kantorovich equivalence [33, Theorem 2.18], without it, equality (a) will become an inequality, which harms finding closed-form $\Theta$. Then it follows that $\frac{t^{\Theta}(x)-\theta_{1}}{\theta_{2}}=\Psi^{-1}(F_{\mu}(x))$, which yields (84). On the other hand, if $\theta_{2}=0$, since $G_{\Theta}(Z)\equiv\theta_{1}\in\mathbb{R}$, we have $F_{\nu^{\Theta}}(y)=1_{\{y\geq\theta_{1}\}}$. Plugging this into (30) directly gives $t^{\Theta}(x)=\theta_{1}$ for all $x\in\mathbb{R}$. This particularly shows that (84) is also satisfied for the case $\theta_{2}=0$. Our target sub-problem becomes (32). As will be explained below, the other sub-problem can be solved in a similar manner.
Together with (8), (86) becomes the constrained optimization problem
$\hbox{subject to}\quad g(\theta_{1},\theta_{2}):=-\theta_{2}\leq 0.$ (87)
@@ Appendix A Proof for Proposition 1
$\theta_{1}\mathbb{E}\left[\Psi^{-1}(F_{\mu}(X))\right]+\theta_{2}\mathbb{E}\left[(\Psi^{-1}(F_{\mu}(X)))^{2}\right]-\mathbb{E}\left[X\Psi^{-1}(F_{\mu}(X))\right]=\lambda/2.$
Note that $F_{\mu}(X)\sim\operatorname{Uniform}[0,1]$, so that the CDF of $\Psi^{-1}(F_{\mu}(X))$ is simply $\Psi$. In other words, $\Psi^{-1}(F_{\mu}(X))$ and $h(Z)$ have identical distribution as (45) in Sec.IV-A. The formulas above thus simplify to
$\theta_{1}+\theta_{2}\mathbb{E}\left[h(Z)\right]-\mathbb{E}[X]=0,$
$\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))\right)\geq 0$ (89)
in (88) following the step of (49) in Sec.IV-A. We therefore separate the proof into two cases from (89).
Case II: $\operatorname{Cov}(X,\Psi^{-1}(F_{\mu}(X)))=0$. As (88) entails $\lambda/2=\theta_{2}\text{Var}(h(Z))$, we have $\lambda=\theta_{2}=0$ (by recalling that either $\lambda=0$ or $\theta_{2}=0$). That is, the KKT condition gives a unique candidate optimizer $(\theta_{1}^{*},\theta_{2}^{*},\lambda^{*})=(\mathbb{E}[X],0,0)$. Since $H_{J}+\lambda H_{g}$ is again given by (91), which is positive definite, we conclude that $(\theta_{1}^{*},\theta_{2}^{*})=(\mathbb{E}[X],0)$ is the unique minimizer of (87).
@@
As (32) is equivalent to (87), Proposition 1 directly shows that $\Theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})\in\mathbb{R}\times\mathbb{R}_{+}$ as in (10) is the unique minimizer of (32). Thanks to (84) and (10),
$\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)$
$=\int_{\mathbb{R}}(x-\mathbb{E}[X])^{2}d\mu(x)-2\theta^{*}_{2}\int_{\mathbb{R}}(x-\mathbb{E}[X])\Big(\Psi^{-1}(F_{\mu}(x))-\mathbb{E}[h(Z)]\Big)d\mu(x)$
$\ \ +(\theta^{*}_{2})^{2}\int_{\mathbb{R}}\Big(\Psi^{-1}(F_{\mu}(x))-\mathbb{E}[h(Z)]\Big)^{2}d\mu(x)$
where the fourth equality follows from the fact that $\Psi^{-1}(F_{\mu}(X))$ and $h(Z)$ have identical distribution, and the last equality is due to the definition of $\theta^{*}_{2}$ in (10). ∎
@@
Suppose that $\mu$ has no atom, the CDF $\Psi$ of $h(Z)$ is continuous and strictly increasing, and $\operatorname{Var}(h(Z))>0$. The sub-problem (56) has a unique minimizer $\Theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})\in\mathbb{R}\times\mathbb{R}_{+}$, defined by
Moreover, the minimum value is